Documenta Mathematica

Journal der Deutschen Mathematiker-Vereinigung Band 1, 1996

Stand vom 3. Dezember 1996

Johan de Jong and Marius van der Put

Étale Cohomology of Rigid Analytic Spaces  1– 56

Nikita A. Karpenko

Order of Torsion in CH^{4} of Quadrics  57– 65

Detlev W. Hoffmann

Twisted Pfister Forms  67–102

Jesús A. de Loera, Serkan Hoşten, Francisco Santos, Bernd Sturmfels The Polytope of All Triangulations

of a Point Configuration  103–119

Friedhelm Waldhausen

On the Construction of the Kan Loop Group  121–126

O. T. Izhboldin

On the Nonexcellence of Field Extensions F (π)/F  127–136

Jürgen Richter-Gebert

Two Interesting Oriented Matroids  137–148

Manfred Knebusch and Digen Zhang

Manis Valuations and Prüfer Extensions I  149–197

Bernd Kawohl

Remarks on Quenching  199–208

Markus Rost

On the Dimension of a Composition Algebra  209–214

Bernold Fiedler

Do Global Attractors Depend on Boundary Conditions?  215–228

A. S. Merkurjev

Maximal Indexes of Tits Algebras  229–243

J. F. Jardine

Boolean Localization, in Practice  245–275

Frans Keune Multirelative K-Theory

and Axioms for the K-Theory of Rings  277–291

Meinolf Geck On the Average Values of the Irreducible Characters of Finite Groups of Lie Type

on Geometric Unipotent Classes  293–317

Markus Rost

Chow Groups with Coefficients  319–393